In a class room which has 30 students, 16 of the students are boys , 23 are ten years old and 6 are girls who are not 10 years old. How many students in the class are ten-year old boys ?
A 8 b 10 c 12 d 15 e 16

Answer by richwmiller(9143)

30-16=14 g
6 g are not 10
14-6=8 g (10)
23-8=15 b(10)
16-1=1 boy not 10
15 b are 10